OpenStudy (anonymous):

(x^2 + 13x + 40) ÷ (x + 8)

OpenStudy (anonymous):

factor out x^2 + 13x + 40 into (x=5)(x+8) so you can cancel something out

OpenStudy (anonymous):

can you write out the steps for me?

OpenStudy (anonymous):

x^2 + 13x + 40 -------------- x + 8 factor out (x+5)(x+8) --------- x+8

OpenStudy (anonymous):

explain pleae!

OpenStudy (anonymous):

\[\frac{a*b}{b} = a\]

OpenStudy (anonymous):

what do you need explaining? how x^2 + 13x + 40 became (x+5)(x+8)?

OpenStudy (anonymous):

yes and the final answer

OpenStudy (anonymous):

Have you studied the quadratic formula, or synthetic division?